OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) please explain! :) thNK YOU!!

OpenStudy (radar):

Step 1. Clear the parens -2x+6>/=5-x-3\[-2x+6\ge-x+2\]Step 3, place unknowns to the left and know values to the right by using legal algebraic operations. \[-x \ge-4\] Step 4. Multiply by -1, and don't forget when multiplying or dividing thru by a negative number (as we are doing here) reverse the sense of the inequality. \[x \le4\]

OpenStudy (radar):

Do you follow with understanding?

OpenStudy (radar):

Hero did not consider the 5 which was in the original equation.

hero (hero):

Yup radar, you got me

OpenStudy (radar):

No, I wasn't trying to get u, I liked the way you avoided having to reverse the sense!.